# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v1

@@ 3.2 Mutual Information
As a metric for correlations, we use MI as in Lin and Tegmark (2017). The MI between two tags is defined by
@@
where $P(x_{0})$ and $P(x_{1})$ are the marginal probabilities of the first and second tags, and $P(x_{0},x_{1})$ is their joint probability. A naive estimate of $I$ can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into Eq. (1). However, this estimation is biased positively due to finite samples (Li, 1990). To reduce this bias, we decompose the MI into
@@
where $S_{0}$, $S_{1}$, and $S_{01}$ are the Shannon entropy, defined by $S_{0}=-\sum_{x_{0}}P(x_{0})\ln P(x_{0})$, $S_{1}=-\sum_{x_{1}}P(x_{1})\ln P(x_{1})$, and $S_{01}=-\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln P(x_{0},x_{1})$. Then we estimate the entropy $S_{0}$ using the method proposed by Grassberger (2003):
@@
where $\psi$ is the digamma function, $n_{x_{0}}$ is the number of samples such that the former tag is $x_{0}$, and $N=\sum_{x_{0}}n_{x_{0}}$ is the total number of samples. $S_{1}$ and $S_{01}$ are estimated in a similar manner. By computing $\hat{S}_{0}$, $\hat{S}_{1}$, and $\hat{S}_{01}$ from the empirical distributions and using Eqs. (2) and (3), we can estimate the MI with a smaller bias.
To evaluate the sufficiency of the sampling, we compute the MI estimate while varying the number of samples, $N_{\text{data}}$. If the estimate converges as $N_{\text{data}}$ increases, this suggests that sampling bias is negligible.
To estimate MI in sequences, we sample $N_{\text{data}}$ pairs of POS nodes separated by a sequential distance $r_{\text{seq}}$, from the dataset. For MI in structures, we sample $N_{\text{data}}$ pairs of nodes at a fixed structural distance $r_{\text{str}}$, where the nodes are either POS or phrasal. We treat ordered pairs as distinct to avoid arbitrarily selecting one of the two. For example, we count pairs (AUX, PP) and (PP, AUX), corresponding to were and than, in Figure 2 as distinct to each other if both pairs are sampled.
@@ 3.4 Context-free Independence Breaking
To examine the validity of PCFGs as models of syntactic structures, it is necessary to quantify the extent to which structures deviate from PCFGs. We employ a metric for this quantification, called context-free independence breaking (CFIB; Nakaishi and Hukushima, 2024).
The essential property of PCFGs, which we refer to as the context-free independence, is that the distribution of states of each node depends only on its parent. Consequently, given the states of two nodes, their children are mutually independent.
Therefore, we can quantify the deviation from PCFGs by the MI between two pairs under the condition that the states of their parents are fixed, as presented in Figure 3. Specifically, provided that the states of two nodes are $x_{0}$ and $x_{1}$, the MI between their children is defined as
@@
Here, $P(y_{0},z_{0}|x_{0},x_{1})$ denotes the conditional probability that the left and right children of the node with label $x_{1}$ are $y_{0}$ and $z_{0}$, respectively. Similarly, $P(y_{1},z_{1}|x_{0},x_{1})$ refers to those for the node with label $x_{1}$. $P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})$ is the conditional joint probability of both pairs of children. By construction, this metric is non-negative for any tree structures and zero for PCFGs.
Figure 3: CFIB $J$ is the MI between the children of two nodes whose states are fixed.
For the estimation of the CFIB, we sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in the previous section. We also distinguish between pairs and their permutations, as in the estimation of MI.
@@ 3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $1\times 10^{2}$ iterations per generation. This process resulted in $13,256,547$ terminated trees. We then apply the same analyses to these generated trees as to parse trees.
@@ 4.1 Mutual Information in Sequences
First, we confirm that the correlation in sequences decays according to a power law. Figure 4 shows the MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$ for different data sizes $N_{\text{data}}$. The estimated values converge as $N_{\text{data}}$ increases, indicating that the estimation bias is sufficiently small. In addition, the decay is better fitted by a power-law function than by an exponential one. This result indicates that the correlation between POS tags within a sentence follows a power law, consistent with previous research. Similar results are also observed for the WikiText corpus, see Appendix D.
@@ 4.2 Mutual Information in Structures
In this subsection, we address question (i), whether the exponential decay of correlation with the structural distance holds in syntactic structures. The estimated MI between POS or phrasal tags as a function of the structural distance is shown in Figure 5. As in the previous subsection, the convergence of the estimated values with increasing $N_{\text{data}}$ indicates that the bias is negligible. Interestingly, model fitting reveals that the correlation decays according to a power law rather than an exponential form with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the previously proposed argument to hold. We have also confirmed that similar results are obtained for the unbinarized case and the WikiText corpus, as detailed in Appendices C and D.
@@ 4.3 Growth of Sequential Distance
The second question is (ii): Does the sequential distance grow exponentially with increasing structural distance? To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 6. We also compute the averages sequential distance for each structural distance, represented by black circles, and fit the results using both exponential and power-law functions. The fitting reveals that the growth follows a power law. However, the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the power law exponent $\beta=0.87$ is smaller than $1$, indicating that the growth is sublinear, i.e., slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a storing branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures. We have also observed similar results for the unbinarized case and for the WikiText corpus in Appendices C and D.
@@ 4.4 Context-free Independence Breaking
Question (iii) asks whether PCFGs are valid models for the hierarchical structures of natural languages. In this section, we measure the CFIB $J$ as the metric of deviations from PCFGs for syntactic structures. In Figure 7, we present the results for fixed nodes being $(x_{0},x_{1})=(\text{NP},\text{NP})$, the most frequent pair in the corpus. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible. Moreover, the CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law with respect to structural distance, suggesting that the context-free independence is significantly broken even at large distances. These findings lead to answering question (iii) in the negative. As shown in Appendix B, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$. In addition, Appendix D shows that the decay of $J$ also follows a power law for the WikiText corpus.
@@ 4.5 PCFG approximation
To deepen our understanding of the results above in comparison with PCFGs, we apply the same analyses to the PCFG constructed to approximate the BLLIP corpus. Figure 8 summarizes the results. The most striking difference between the parse trees and the PCFG appears in the CFIB metric, shown in Figure 8 (d). The estimated value of this metric does not saturate but continues to decrease even for large data sizes. This behavior is consistent with the fact that CFIB is identically zero for PCFGs by definition.
For the other statistical quantities, the differences are more moderate. As shown in Figure 8 (a), the MI in sequences decays according to a power law at short sequential distances, $r_{\text{seq}}\lesssim 10$, while for large sequential distances, $r_{\text{seq}}\gtrsim 10$, it increases substantially, in contrast to the trend in parse trees. The MI in structures, as shown in Figure 8 (b), exhibits a power-law decay as in parse trees. In Figure 8 (c), the average sequential distance grows more slowly than exponentially and even than linearly, indicating that the generated trees are strongly biased. While this behavior is qualitatively similar to that in syntactic structures, the estimated exponent $\beta=0.45$ significantly differs from that for the parse trees, where $\beta=0.87$.
@@
Although these findings may suggest that the PCFG partially captures the statistical properties of parse trees, the mechanism underlying the power-law decay of correlation in sequences clearly differs from the one proposed in Lin and Tegmark (2017), in which the decay of MI in structures and the growth of sequential distance are exponential. Importantly, our finding that the MI in structures decays according to a power law in the PCFG suggests that the distance within a sentence is too short for the MI to converge to its asymptotic exponential decay.
This indicates that the proposed argument is difficult to apply to short sequences such as single sentences. In such cases, it becomes impossible to distinguish whether the process underlying the sequences is Markov-like or PCFG-like. This limitation also arises when analyzing child languages and animal languages, because available datasets do not contain sequences of sufficient length. Indeed, previous studies have analyzed sequences shorter than $\sim 10^{2}$ phones, syllables, words, or action elements (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).

### v2

@@ 3.2 Mutual Information
As a metric for correlations, we use the mutual information (MI) as in Lin and Tegmark (2017). The MI between two tags is defined by
@@
where $P(x_{0})$ and $P(x_{1})$ are the marginal probabilities of the first and second tags, and $P(x_{0},x_{1})$ is their joint probability. The MI can be decomposed into
@@
where $S_{0}$, $S_{1}$, and $S_{01}$ are the Shannon entropy, defined by $S_{0}=-\sum_{x_{0}}P(x_{0})\ln P(x_{0})$, $S_{1}=-\sum_{x_{1}}P(x_{1})\ln P(x_{1})$, and $S_{01}=-\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln P(x_{0},x_{1})$. By estimating $S_{0}$, $S_{1}$, and $S_{01}$, we can estimate the MI.
A naive estimate can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into $P(x_{0})$, $P(x_{1})$, and $P(x_{0},x_{1})$ in the definitions of entropy. In this case, the estimates of entropy and MI are biased due to finite samples (Arora et al., 2022; Li, 1990).
To reduce this bias, we employ the more sophisticated estimator proposed by Grassberger (2003). In this method, the entropy $S_{0}$ is estimated by
@@
where $\psi$ is the digamma function, $n_{x_{0}}$ is the number of samples such that the former tag is $x_{0}$, and $N=\sum_{x_{0}}n_{x_{0}}$ is the total number of samples. $S_{1}$ and $S_{01}$ are estimated in a similar manner. Grassberger (2003) has proven that the bias of this estimator rapidly converges to zero as the number of samples, $N$, increases. Therefore, if the estimated value saturates with respect to $N$, the finite-sample bias should be negligible.
To estimate the MI in sequences, we sample $N_{\text{data}}$ pairs of POS nodes separated by a sequential distance $r_{\text{seq}}$, from the dataset. For the MI in structures, we sample $N_{\text{data}}$ pairs of nodes at a fixed structural distance $r_{\text{str}}$, where the nodes are either POS or phrasal. We treat ordered pairs as distinct to avoid arbitrarily selecting one of the two. For example, we count pairs (AUX, PP) and (PP, AUX), corresponding to were and than, in Figure 2 as distinct to each other if both pairs are sampled.
In the following analyses, we compute the MI estimates while varying the number of samples, $N_{\text{data}}$. In our results, the estimates saturate as $N_{\text{data}}$ increases, indicating that the bias is sufficiently small.
To further support our claim, we also report the results obtained with six alternative entropy estimators Madow (1948); Miller (1955); Zahl (1977); Horvitz and Thompson (1952); Chao and Shen (2003); Wolpert and Wolf (1995); Nemenman et al. (2001) in Appendix C. The results are consistent with those reported in the main text, except for the estimators that were previously reported to exhibit larger errors (Arora et al., 2022).
@@ 3.4 Context-free Independence Breaking
To address question (iii), it is necessary to quantify the extent to which syntactic structures deviate from PCFGs. We employ a metric for this quantification, called context-free independence breaking (CFIB; Nakaishi and Hukushima, 2024).
The essential property of PCFGs, which we refer to as the context-free independence, is that the distribution of categories of each node depends only on its parent. Consequently, given the categories of two nodes, their children are mutually independent.
Figure 3: CFIB $J$ is the MI between the children of two nodes whose categories are fixed.
Therefore, we can quantify the deviation of the distribution of trees from PCFGs by the MI between two pairs under the condition that the categories of their parents are fixed, as presented in Figure 3. Specifically, provided that the categories of two nodes are $x_{0}$ and $x_{1}$, the CFIB is defined as
@@
Here, $P(y_{0},z_{0}|x_{0},x_{1})$ denotes the conditional probability that the left and right children of the node with label $x_{1}$ are $y_{0}$ and $z_{0}$, respectively. Similarly, $P(y_{1},z_{1}|x_{0},x_{1})$ refers to those for the node with label $x_{1}$. $P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})$ is the conditional joint probability of both pairs of children. By definition, the CFIB is non-negative for any tree structures and zero for trees generated by PCFGs.
The CFIB is defined for the distribution of trees, as is the MI. We can estimate this metric from the dataset of trees, similarly to the estimation of the MI. We sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in the previous section. We also distinguish between pairs and their permutations, as in the estimation of MI.
We verify the accuracy of the estimates for the CFIB by varying $N_{\text{data}}$, as we did for the MI. Appendix C also presents the results obtained with alternative estimators, further supporting the validity of our estimates.
@@ 3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of natural language parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $1\times 10^{2}$ iterations per generation.This procedure yields approximately 13 million terminated trees. We then apply the same analyses to these generated trees as to parse trees.
@@ 3.6 Fitting
3.7 Synthetic Data
To demonstrate that our method can reliably estimate MI and distinguish between exponential and power-law decays, we apply it to mathematical models in which MI can be computed exactly. We generate random variables $X$ and $Y$ with joint probabilities defined as $P(X=0,Y=0)=P(X=1,Y=1)=(1+\delta)/4$ and $P(X=0,Y=1)=P(X=1,Y=0)=(1-\delta)/4$. The MI between $X$ and $Y$ is $I=\frac{1+\delta}{2}\ln(1+\delta)+\frac{1-\delta}{2}\ln(1-\delta)$, which is $\delta^{2}/2+\mathcal{O}(\delta^{3})$ as $\delta\to 0$. If $\delta=\exp(-\lambda r/2)$, the MI decays asymptotically as $I\sim\exp(-\lambda r)$. If $\delta=r^{-\alpha/2}$, the decay follows a power law, $I\sim r^{-\alpha}$. We refer to these as the exponential and power-law models, respectively.
Figure 4: Estimation and fitting of the MI for (a) the exponential model with $\lambda=0.1$ and (b) the power-law model with $\alpha=2$. Fitting was performed for $N_{\text{data}}=8\times 10^{6}$.
We sample $X$ and $Y$ from the exponential model with $\lambda=0.1$ and from the power-law model with $\alpha=2$. We then estimate the MI and fit the resulting values. The results are shown in Figure 4. In both cases, the estimates converge to the exact values as the sample size $N_{\text{data}}$ increases. In addition, $\chi^{2}_{\nu}$ is significantly smaller when the fitting function matches the model. Our method successfully discriminates between the exponential and power-law models.
@@ 4.1 Mutual Information in Sequences
First, we confirm that the correlation in sequences decays according to a power law. Figure 5 shows the MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$ for different data sizes $N_{\text{data}}$. The estimated values converge as $N_{\text{data}}$ increases, indicating that the estimation bias is sufficiently small.
The decay of $I_{\text{POS}}$ is better fitted by a power-law function ($\chi^{2}_{\nu}=2.6\times 10^{-1}$) than by an exponential one ($\chi^{2}_{\nu}=9.8\times 10^{-1}$). This result indicates that the correlation between POS tags follows a power law, consistent with previous research.
@@ 4.2 Mutual Information in Structures
We now address question (i): whether the exponential decay of correlation with the structural distance holds in natural language syntactic structures. The estimated MI between POS or phrasal tags as a function of the structural distance is shown in Figure 6. As in the previous section, the convergence of the estimated values with increasing $N_{\text{data}}$ indicates that the bias is negligible.
Model fitting reveals that the correlation decays according to a power law ($\chi^{2}_{\nu}=1.1\times 10^{-1}$) rather than an exponential form ($\chi^{2}_{\nu}=1.3$) with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the previously proposed argument to hold.
@@ 4.3 Growth of Sequential Distance
Question (ii) asks whether the sequential distance grows exponentially with increasing structural distance. To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 7. We also compute the average sequential distance for each structural distance, represented by black circles, and fit the results using both exponential and power-law functions.
The fitting reveals that the growth follows a power law ($\chi^{2}_{\nu}=1.4\times 10$) rather than an exponential form ($\chi^{2}_{\nu}=1.4\times 10^{21}$), although the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the power law exponent $\beta=0.72$ is smaller than $1$, indicating that the growth is slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a strong branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures.
@@ 4.4 Context-free Independence Breaking
Question (iii) concerns whether the statistical properties of syntactic structures deviate significantly from those of PCFGs. We measure the CFIB $J$ as a metric of this deviation. In Figure 8, we present the results for fixed nodes being $(x_{0},x_{1})=(\text{NP},\text{NP})$, the most frequent pair in the corpus. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible.
The CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law ($\chi^{2}_{\nu}=8.1\times 10^{-2}$) rather than an exponential form ($\chi^{2}_{\nu}=7.1\times 10^{-1}$) with respect to structural distance, suggesting that the context-free independence is significantly broken even at large distances. These results indicate that the answer to question (iii) is negative.
This finding is also consistent with previous studies that developed PCFG-based parsers achieving high performance (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). These studies augmented PCFGs to relax the context-free independence implicitly or explicitly.
@@
As shown in Appendix D, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$, suggesting that the observed deviation from PCFGs is not specific to a particular syntactic category pair.
4.5 Other Corpora
We examined whether the three assumptions in the argument of Lin and Tegmark (2017) hold for BLLIP. The results show that all of these assumptions are violated. The argument cannot account for the statistical properties of parse trees in this corpus. To test whether this conclusion is specific to BLLIP, we conduct the same analyses on WikiText and NPCMJ.
Figure 9: Statistical properties of syntactic structures in WikiText and NPCMJ. (a1–4) Results for WikiText: (a1) $I_{\text{POS}}$ as a function of $r_{\text{seq}}$; (a2) $I_{\text{tag}}$ as a function of $r_{\text{str}}$; (a3) relationship between $r_{\text{seq}}$ and $r_{\text{str}}$; and (a4) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ as a function of $r_{\text{str}}$. Fitting was performed with $N_{\text{data}}=2.56\times 10^{6}$. (b1–b4) Corresponding results for NPCMJ. Fitting was performed with $N_{\text{data}}=6.4\times 10^{5}$ for $I_{\text{POS}}$ and $I_{\text{tag}}$, and with $N_{\text{data}}=4.0\times 10^{4}$ for $J$.
Figure 9 summarizes the results. For both corpora, the statistical properties are similar to those observed for BLLIP: the MI decays according to a power law in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These behaviors are supported by the chi-square statistics, which are smaller for power-law fitting than for exponential fitting across all statistical quantities.
The results for these corpora also exhibit slight differences from those for BLLIP. $I_{\text{POS}}$ and $I_{\text{tag}}$ for WikiText, as well as $I_{\text{POS}}$ for NPCMJ, increase at large distances. This effect may be attributable to structural differences in unusually long sentences or to potential parsing errors in such cases. However, we cannot determine the significance of this behavior because the datasets contain too few long sentences.
Nonetheless, the results clearly support that the assumptions proposed by Lin and Tegmark (2017) are violated not only for BLLIP, but also for another English corpus, WikiText, as well as a Japanese treebank, NPCMJ. These results imply the robustness of our conclusion, given that Japanese is typologically distinct from English, exhibiting head-final, left-branching syntactic structures.
@@ 4.6 PCFG Approximation
To further elucidate the difference between syntactic structures and PCFGs, we apply the same analyses to the PCFG constructed using maximum likelihood estimation on BLLIP. The results are summarized in Figure 10.
Mutual Information in Sequences
Figure 10 (a) shows the MI in sequences. At short sequential distances, $r_{\text{seq}}\lesssim 10$, the MI decays according to a power law, although the decay rate does not clearly match that observed in parse trees. At larger distances, $r_{\text{seq}}\gtrsim 10$, the MI increases substantially, in contrast to the trend in parse trees.
Mutual Information in Structures
The MI in structures, shown in Figure 10 (b), exhibits a power-law decay similar to that in syntactic structures, although the decay rate differs. This result does not contradict the properties of PCFGs proven by Lin and Tegmark (2017). While the MI in structures generated by PCFGs asymptotically follows an exponential decay at sufficiently large distances, the MI can exhibit different behavior at shorter distances. The non-exponential behavior of $I_{\text{tag}}$ suggests that for individual sentences, the structural distances are too short for the asymptotic behavior to emerge.
If the generation process continued for more steps, producing sufficiently large trees, the MI in Figure 10 (b) would converge to an exponential decay at large structural distances. However, such a setup cannot be justified because the generated trees would be much larger than typical parse trees. Even if we assume this setup, the MI in sequences would decay more rapidly than a power law as long as the trees are sufficiently biased so that the sequential distances grow more slowly than exponentially. The mechanism of the power-law decay of correlation observed in natural languages remains unexplained.
The non-exponential behavior of $I_{\text{tag}}$ in the PCFG has another important implication. As discussed in Section 2.2, the argument by Lin and Tegmark (2017) can, in principle, be extended to hierarchical structures beyond syntax. Following this perspective, several studies have applied the argument to domains other than texts and speeches by human adults. These studies reported the power-law decay of correlation in child speeches and animal signals, interpreting it as evidence for the existence of underlying hierarchical structures in these domains (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, the argument by Lin and Tegmark (2017) relies on the asymptotic behavior of PCFGs, which does not appear when the observed sequences are too short. In child speeches and animal signals, the sequences are typically shorter than sentences produced by human adults. In such cases, it is difficult to interpret the power-law decay of correlation as evidence for hierarchical structures.
Growth of Sequential Distance
In Figure 10 (c), the average sequential distance grows more slowly than exponentially and even than linearly, indicating that the generated trees are strongly biased. While this behavior is qualitatively similar to that in parse trees, the estimated exponent $\beta=0.41$ significantly differs from that for the parse trees, where $\beta=0.72$.
Context-free Independence Breaking
The most striking difference between parse trees and trees generated by the PCFG appears in the CFIB metric. The true value of this metric is zero for PCFGs by definition. If the estimates converge to this true value with increasing data size, the corresponding values in the log-scale plot continue to decrease, approaching -$\infty$, since the logarithm of zero is -$\infty$.
In Figure 10 (d), the estimates of CFIB is much smaller than those for syntactic structures in Figure 8. Furthermore, the estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is the expected behavior. This result is in contrast to that for syntactic structures in natural languages.